OpenStudy (anonymous):

A company with loud machinery needs to cut its sound intensity to 44% of its original level. By how many decibels would the loudness be reduced? Use the formula L=10log I/I(o). Round to the nearest hundredth.

OpenStudy (anonymous):

3.57 decibels 1.58 decibels 2.52 decibels 1.93 decibels

OpenStudy (anonymous):

If the sound intensity is reduced forty four percent from what we had before, we know that\[L_{new}=10\log(\frac{.56I}{I_0})=10(\log(\frac{I}{I_0})+\log(.56))\]\[=L_{old}+10\log(.56) \approx L_{old}-2.52\]

OpenStudy (anonymous):

@uri how did they DO THIS?